preliminary draft, 15 May 1995 17:19  preliminary draft, 15 May 1995 17:19 TUGboat, Volume 0 (2001), No. 0  preliminary draft, 15 May 1995 17:19  1001

First report from the DVI driver standards committee

Don Hosek

The TEX Users Group DVI driver standards committee has been working on the development of standards for device drivers since the fall of 1988. This article is a first report on our status to the membership of TEX Users Group.

At the time of this writing, we are in the midst of discussion of \special standards for device drivers. By the TEX Users Group meeting this August, we should have a preliminary report on this topic available for distribution to all interested parties. We welcome all input from members of the TEX community; if you have any suggestions, comments, etc. regarding the issue of \special handling, we would appreciate it if you could send these to Robert McGaffey (internet:

Mcgaffey%Orn.Mfenet@Nmfecc.Arpa) for distribu-

tion to the members of the committee.

The members of the committee are: Robert McGaffey, chair, Oak Ridge National Laboratory; David P. Babcock, Hewlett-Packard; Elizabeth Barnhart, TV Guide; Stephan v. Bechtolsheim, Integrated Computer Software Inc.; Nelson Beebe, University of Utah; Jackie Damrau, University of New Mexico; Donald Goldhammer, University of Chicago; Don Hosek, University of Illinois at Chicago; David Ness, TV Guide; Thomas J. Reid, Texas A&M University; David Rodgers, Arbortext, Inc.; Brian Skidmore, Addison-Wesley Publishing Co.; Glenn Vanderburg, Texas A&M University; and Ralph Youngen, American Mathematical Society.